OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function y = (cosx)^x

OpenStudy (anonymous):

take the log both sides. lny=xln(cosx)

OpenStudy (anonymous):

and take derivative both sides

OpenStudy (anonymous):

y prime/y= (ln(cosx))+x(1/cosx)(-sinx) and isolate y prime and you are done

OpenStudy (anonymous):

\[\int\limits \text{Cos}[x]^x (\text{Log}[\text{Cos}[x]]-x \text{Tan}[x])dx = \text{Cos}[x]^x \]

OpenStudy (anonymous):

put ln on both sides to get \[\ln y =x \ln cosx\] then differentiate to get \[1/y =1\times -sinx\]

OpenStudy (anonymous):

also you can use formula for logarithmic differentiation... d/dx(u^v)=u^v*d/dx(v*log u )